# Beyond The Rainbow: High Performance Deep Reinforcement Learning On A Desktop PC

### v1

@@ 3.2 Hyperparameters
For many years, RL algorithms have used a discount rate of $0.99$, however, when reaching high performance, lower discount rates alter the optimal policy, causing even optimally performing agents to not collect the maximum cumulative rewards. To prevent this, we follow MuZero Reanalyse (Schrittwieser et al., 2021) using $\gamma=0.997$. For our Prioritized Experience Replay, we use the lower value of $\alpha=0.2$, the parameter used to determine sample priority, recommended by Toromanoff et al. (2019) when using IQN. Lastly, many previous experiments used only noisy networks or $\epsilon$-greedy exploration, however, we opt to use both until 100M frames, then set $\epsilon$ to zero, effectively disabling it. We elaborate on this decision in Appendix G.
@@ 4 Evaluation
To assess BTR, we test it on two standard RL benchmarks, Atari (Bellemare et al., 2013) and Procgen (Cobbe et al., 2020) in Section 4.1. Secondly, we train BTR agents for three modern games (Super Mario Galaxy, Mario Kart Wii, and Mortal Combat) with complex 3D graphics and physics in Section 4.2, never shown to be trainable with RL previously.
@@ 4.1 Atari and Procgen Performance
We evaluate BTR on the Atari-60 benchmark following (Machado et al., 2018) and without life information (see Appendix J for the impact), evaluating every million frames on 100 episodes. Figure 1 plots BTR against Rainbow DQN and DQN, achieving an IQM of 7.4 compared to Rainbow DQN’s 2.7 and DQN’s 0.9. In comparison to human expert performance, BTR equals or exceeds them in 52 of 60 and Rainbow DQN in 57. Importantly, we find that BTR appears to continue increasing performance beyond 200 million frames, indicating that higher performance is still possible with more time and data. Results tables and graphs can be found in Appendices A and B respectively.
Figure 2: IQM min-max Performance of BTR compared to Rainbow DQN + Impala (width x4) Cobbe et al. (2020) after 200M frames on all 16 Procgen Environments.
To further confirm BTR’s performance, we benchmark on Procgen (Cobbe et al., 2020), a procedurally generated set of environments aiming to prevent overfitting to specific tasks, a prevalent problem in RL (Justesen et al., 2018; Juliani et al., 2019). The results are shown in Figure 2 with individual games in Appendix B. BTR is able to exceed Rainbow DQN + Impala’s performance, despite using significantly fewer convolutional filters (which Cobbe et al. (2020) found to significantly improve performance) and using 8 hours of walltime compared to 41. These results demonstrate BTR’s general learning capability across a wide range of standard RL benchmarks.
@@ 4.2 Applying BTR to Modern Games
To demonstrate BTR’s capabilities beyond standard RL benchmarks, we utilised Dolphin (Dolphin-Emulator, 2024), a Nintendo Wii emulator, to train agents for a range of modern 3D games: Super Mario Galaxy, Mario Kart Wii and Mortal Combat. Using a desktop PC, we were able to train the agent to complete some of the most difficult tasks within each game. Namely, the final level in Super Mario Galaxy, Rainbow Road (a notoriously difficult track in Mario Kart Wii) and defeating all opponents in Mortal Kombat Endurance mode (for details about the environments and setup, see Appendix K). For this, BTR required minimal adjustments: first, to input image resolution, 140x114 (from Atari’s 84x84) due to the game’s higher resolution and aspect ratio, and second, to reduce the number of vectorized environments to 4 as a result of the games’ memory and CPU requirements.
Figure 3: BTR being used to play Super Mario Galaxy (final level), Mario Kart Wii (Rainbow Road) and Mortal Kombat: Armageddon (Endurance Mode) respectively.
For all the games, BTR was able to solve the game level, including consistently finishing in first place in Mario Kart. We provide videos of our agent playing all three Wii games, in addition to the games in the Atari-5 benchmark  [^5] https://www.youtube.com/playlist?list=PL4geUsKi0NN-sjbuZP_fU28AmAPQunLoI.
@@ 5 Analysis
Given BTR’s performance demonstrated in Section 4, in this Section, we ablate each component to evaluate their performance impact (Section 5.1). Using the ablated agents, we then measure numerous attributes during and after training to assess each component’s impact (Section 5.2).
@@ 5.1 Ablations Studies
BTR amalgamates independently evaluated components into a single algorithm. To understand and verify each’s contribution, Figures 4 and B1 plot BTR’s performance without each component on the Atari-5 benchmark. [^6] Due to computational resources required to evaluate each component on 60 environments, Aitchison et al. (2023) proposes a subset of 5 games that closely correlate with the performance across all 60.
Figure 4: Individual game performance of BTR on Atari-5 with individual components removed averaged over 3 seeds. Shaded areas show 1 standard deviation.
We find that Impala had the largest effect on performance, with the other components generally causing a less significant effect on final performance. However, when BTR’s performance is compared before 200 million frames, we find Munchausen and Spectral Normalisation provide significant performance improvements ($+24\%$ and $+25\%$ at 40M frames, and $+13\%$ and $+35\%$ at 120M frames). We compare the performance of components at different stages of training in Appendix E.
For vectorization and maxpooling, while their inclusion reduces performance, we find their secondary effects crucial to keep BTR computationally accessible. Omitting vectorization increases walltime by 328% (Figure 5) by processing environment steps in parallel and taking fewer gradient steps (781,000 compared to Rainbow DQN’s 12.5 million). [^7] Another consequence of removing vectorization is using smaller batches, which Obando Ceron et al. (2024) finds improves exploration, possibly explaining our results found in Qbert. We find maxpooling makes the agent more robust to noise as discussed in Section 5.2, and decreases the model’s parameters by 77%.
Figure 5: Walltime of BTR on a desktop PC with components removed, compared with Hessel et al. (2018) and Schmidt & Schmied (2021). For hardware details, see Appendix H.
@@ 5.2 What are the effects of BTR’s components?
To help interpret the results in Section 5.1, we measure seven different attributes of the agent either during or after training: action gaps and action swaps, linked to causing approximation errors (Bellemare et al., 2016); policy churn, which can cause excessive off-policyness (Schaul et al., 2022); score with additional noise indicating robustness and weight norm and SRank (Sokar et al., 2023), correlated with loss of plasticity.
@@
For why Impala contributes to performance so strongly, Figure 6 shows that with Impala, BTR has fewer dormant neurons and higher weight matrix norms. Despite this, without BTR’s other components, Impala exhibits a notable drawback, learning a highly noisy and unstable policy. Table 2, demonstrates that without IQN and Munchausen the agent experiences very low action gaps (absolute Q-value difference between the highest two valued actions), causing the agent to swap its argmax action almost every other step. This is likely to result in approximation errors altering the policy and causing a high degree of off-policyness in the replay buffer. This is particularly detrimental in games requiring fine-grained control, such as Phoenix where the agent needs to narrowly dodge many projectiles, reflected in BTR’s performance without these components.
@@
Furthermore, we find that spectral normalization and maxpooling are useful in dealing with noisy environments. To test this, we evaluate the performance of BTR’s ablations when taking different quantities of $\epsilon$-actions, and find these two components prevent performance from dropping substantially 2. Lastly, we find Munchausen and IQN to have a significant impact on Policy Churn (Schaul et al., 2022), with Munchausen reducing it by 5.6% and IQN increasing it by 2.4%. As a result, when these components are used together, they appear to reach a level of policy churn which does not harm learning and potentially provides some exploratory benefits.
Table 2: Comparison of policy churn, action gaps, actions swaps and evaluation performance with different quantities of $\epsilon$-actions and color jitter (both only applied for evaluation). All measurements use the final agent, trained on 200 million frames, for Atari Phoenix. Action Gap is the average absolute Q-value difference between the highest two valued actions. % Actions Swap is the percentage of times the agent’s argmax action has changed from the last timestep. Policy churn is the percentage of states which the agent’s argmax action has changed on after a single gradient step. Color jitter applies a random 10% change to the brightness, saturation and hue of each frame.
Category BTR w/o Munchausen w/o IQN w/o SN w/o Impala w/o Maxpool
Action Gap 0.276 0.060 0.154 0.280 0.403 0.282
% Action Swaps 33.9% 47.0% 45.2% 42.2% 30.9% 39.7%
Policy Churn 2.9% 8.5% 0.5% 2.1% 3.5% 2.8%
Score ColorJitter 196k 90k 97k 167k 5k 149k
Score $\epsilon=0.03$ 96k 53k 49k 86k 5k 72k
Score $\epsilon=0.01$ 211k 88k 89k 173k 5k 138k
Score $\epsilon=0$ 351k 211k 196k 366k 5k 500k
6 Related Work
The most similar work to BTR, developing a computationally-limited non-distributed RL algorithm, is “Fast and Efficient Rainbow” (Schmidt & Schmied, 2021). They optimised Rainbow DQN to maximise performance for 10 million frames through parallelizing the environments and dropping C51 along with hyperparameter optimisations. This differs from our goals of producing an algorithm that scales across training regimes (up to 200 million frames) and domains (Atari, Procgen, Super Mario Galaxy, Mario Kart and Mortal Combat), resulting in different design decisions.
For less computation-limited approaches, Ape-X (Horgan et al., 2018) was the first to explore highly distributed training, allowing agents to be trained on a billion frames in 120 hours through using   100 CPUs. Following this, Kapturowski et al. (2018) proposed R2D2 using a recurrent neural network, increasing sample efficiency but slowing down gradient updates by 38%. Agent57 (Badia et al., 2020a) was the first RL agent to achieve superhuman performance across 57 Atari games, though required 90 billion frames. MEME (Kapturowski et al., 2022), Agent57’s successor, focused on achieving superhuman performance within the standard 200 million frames limit, achieved by used a significantly higher replay ratio and larger network architecture. Most recently, Dreamer-v3 (Hafner et al., 2023) used a 200 million parameter model requiring over a week of training, achieving similar results as MEME. We detail some of the key differences between BTR, MEME and Dreamer-v3 in Table 3. While these approaches perform equally to or better than BTR, all are inaccessible to smaller research labs or hobbyists due to their required computational resources and walltime. Therefore, while these algorithms have important research value demonstrating the possible performance of RL agents, performative algorithms with a lower cost of entry, like BTR, are a necessary component for RL to become widely applicable and accessible.
Figure 6: Plot showing L2 norm of network weights, SRank $\delta=0.01$ (Kumar et al., 2020) and % of dormant neurons (dormant defined as $<$ 0.1 (Sokar et al., 2023), for details see Appendix F).
@@ 6 Related Work
7 Conclusion and Future work
@@
We acknowledge that there are many more promising improvements we were not able to include in BTR, leaving room for more future work in a few years to create even stronger integrated agents. For example, BTR does not add an explicitly exploration component, resulting in it struggling in hard-exploration tasks such as Montezuma’s Revenge; therefore, mechanisms used in Never Give Up (Badia et al., 2020b) or other components may prove useful. Section 5.1 found that the neural network’s core architecture, Impala, had the largest impact on performance, an area we believe is generally underappreciated in RL. Previous work (Kapturowski et al., 2018) has incorporated recurrent models enhancing performance, though we are uncertain how this can be incorporated into BTR without affecting its computational accessibility.
8 Ethics and Reproducibility Statements
Our work does not involve human subjects or methodologies with direct ethical concerns such as discrimination, bias, or privacy violations. Additionally, we have no conflicts of interest, sponsorship issues, or violations of legal or research integrity were present during the development of this research. However, we acknowledge that by improving the accessibility and performance of Reinforcement Learning (RL), our contributions may inadvertently provide more powerful tools to malicious actors, thus, we urge the research community to remain vigilant regarding these issues.
To ensure reproducibility, we provide a detailed background of the work we build upon and clearly explain all changes made to the base algorithm. Furthermore, Appendix C.2 provides all relevant environmental and algorithmic hyperparameters needed to reproduce our work. Additionally, we provide clarity about often misunderstood terms (Appendix C.3), a detailed architecture diagram (Appendix D) and the exact hardware we tested our algorithms on (Appendix H). Most importantly, we provide BTR’s code within the supplementary material. Lastly, we provide many details regarding the Wii games tested BTR on, including the minor changes from BTR, how the environment was setup and the reward functions used.

### v2

@@ 3.2 Hyperparameters
For many years, RL algorithms have used a discount rate of $0.99$, however, when reaching high performance, lower discount rates alter the optimal policy, causing even optimally performing agents to not collect the maximum cumulative rewards. To prevent this, we follow MuZero Reanalyse (Schrittwieser et al., 2021) using $\gamma=0.997$. For our Prioritized Experience Replay, we use the lower value of $\alpha=0.2$, the parameter used to determine sample priority, recommended by Toromanoff et al. (2019) when using IQN. Lastly, many previous experiments used only noisy networks or $\epsilon$-greedy exploration, however, we opt to use both until 100M frames, then set $\epsilon$ to zero, effectively disabling it. We elaborate on this decision in Appendix F.
@@ 4 Evaluation
To assess BTR, we test on two standard RL benchmarks, Atari (Bellemare et al., 2013) and Procgen (Cobbe et al., 2020) in Section 4.1. Secondly, we train agents for three modern games (Super Mario Galaxy, Mario Kart Wii, and Mortal Kombat) with complex 3D graphics and physics in Section 4.2, never before shown to be trainable with RL.
Figure 2: Box plot performance of BTR (4 seeds) against other popular algorithms such as MEME (Kapturowski et al., 2023), Dreamer v3 (Hafner et al., 2023), Bigger, Better, Faster (BBF) (Schwarzer et al., 2023) and EfficientZero-v2 (EZV2) (Wang et al., 2024). Brackets show the number of frames the algorithms use, the number of walltime hours and the hardware used respectively. Shaded areas show 95% confidence intervals. Top: Atari 55 game benchmark - we used the overlapping games 55 between the popular Atari-57 benchmark, and the 60 games used in RLiable (Agarwal et al., 2021). Bottom: Atari-26 benchmark, commonly used for testing sample-efficient algorithms.
Figure 3: BTR compared to Rainbow DQN + Impala (width x4) (Cobbe et al., 2020) after 200M frames on the Procgen benchmark. Shaded areas show 95% CIs, with results averaged over 5 seeds.
Figure 4: BTR being used to play Super Mario Galaxy (final level), Mario Kart Wii (Rainbow Road) and Mortal Kombat: Armageddon (Endurance Mode) respectively. Consistent completion is defined as over 90%.
@@ 4.1 Atari and Procgen Performance
We evaluate BTR on the Atari-60 benchmark following (Machado et al., 2018) and without life information (see Appendix I for the impact), evaluating every million frames on 100 episodes. Figure 1 plots BTR against DQN, Rainbow DQN and Dreamer-v3, showing BTR’s competitive performance despite using significantly less walltime. Figure 2 shows a box plot comparison of final performance. In comparison to human expert performance, BTR equals or exceeds them in 52 of 60. Importantly, we find that BTR appears to continue increasing performance beyond 200 million frames, indicating that higher performance is still possible with more time and data. Results tables and graphs can be found in Appendices A and B, respectively.
Figure 5: BTR’s human-normalized scores without different components, with shaded areas showing 95% bootstrapped confidence intervals averaged over 4 seeds. Left: Predicted Atari-57 median score using the regression procedure defined in Aitchison et al. (2023). However, we find the prediction does not match the true median (see Appendix K). Right: Interquartile mean across the 5 games. For individual game graphs and additional ablations, see Appendices B and C.
To further confirm BTR’s performance, we benchmark on Procgen (Cobbe et al., 2020), a procedurally generated set of environments aiming to prevent overfitting to specific tasks, a prevalent problem in RL (Justesen et al., 2018; Juliani et al., 2019). The results are shown in Figure 3 with individual games in Appendix B. BTR is able to exceed Rainbow DQN + Impala’s performance, despite using significantly fewer convolutional filters (which Cobbe et al. (2020) found to significantly improve performance) and using 8 hours of walltime compared to 41. While BTR provides an improvement over Rainbow DQN in Procgen, we did not target procedurally generated environments thus it does not currently compete with the state-of-the-art (Cobbe et al., 2021; Hafner et al., 2023). There are numerous ways performance can be improved (Jesson & Jiang, 2024; Cobbe et al., 2020) which we leave to future work.
@@ 4.2 Applying BTR to Modern Games
To demonstrate BTR’s capabilities beyond standard RL benchmarks, we utilized Dolphin (Dolphin-Emulator, 2024), a Nintendo Wii emulator, to train agents for a range of modern 3D games: Super Mario Galaxy, Mario Kart Wii and Mortal Kombat. Using a desktop PC, we were able to train the agent to complete some of the most difficult tasks within each game. Namely, the final level in Super Mario Galaxy, Rainbow Road (a notoriously difficult track in Mario Kart Wii), and defeating all opponents in Mortal Kombat Endurance mode. For details about the environments and setup, see Appendix J. To achieve this, BTR required minimal adjustments: changing the input image resolution to 140x114 (from Atari’s 84x84) due to the game’s higher resolution and aspect ratio, and to reduce the number of vectorized environments to 4 due to the games’ memory and CPU requirements.
BTR was able to solve all three games, including consistently finishing first place in Mario Kart. In contrast, Rainbow DQN’s performance plateaued before completing any of the games. We provide videos of our agent playing all three Wii games and all games in the Atari-5 benchmark  [^5] https://www.youtube.com/playlist?list=PL4geUsKi0NN-sjbuZP_fU28AmAPQunLoI.
@@ 5 Analysis
Given BTR’s performance demonstrated in Section 4, in this section, we ablate each component to evaluate their performance impact (Section 5.1). Using the ablated agents, we measure numerous attributes during and after training to assess each component’s impact (Section 5.2).
@@ 5.1 Ablations Studies
BTR amalgamates independently evaluated components into a single algorithm. To understand and verify each component’s contribution, Figure 5 plots BTR’s performance without each component on the Atari-5 benchmark. [^6] Due to the resources required to evaluate on all environments, Aitchison et al. (2023) proposes a subset of 5 games that closely correlate with the performance across all of them.
We find that Impala had the largest effect on performance (+142% IQM), with the other components generally causing a less significant effect. Despite this, simply using Rainbow with Impala does not produce similar results (6.3 IQM compared to 7.7 on Atari-5). Munchausen and IQN have a strong impact on environments requiring fine-grained control such as Phoenix, as explored in Section 5.2.
For vectorization and maxpooling, while their inclusion reduces performance, we find their secondary effects crucial to keep BTR computationally accessible. Omitting vectorization increases walltime by 328% (Figure 6) by processing environment steps in parallel and taking fewer gradient steps (781,000 compared to Rainbow DQN’s 12.5 million). [^7] A result of removing vectorization is using smaller batches, which Obando Ceron et al. (2024) finds improves exploration. We find that maxpooling decreases the model’s parameters by 77%, and makes using wider convolutional layers possible without causing the total number of parameters to increase drastically.
Figure 6: Walltime of BTR on a desktop PC with components removed, compared with Hessel et al. (2018) and Schmidt & Schmied (2021). For hardware details, see Appendix G.
Table 2: Comparison of policy churn, action gaps, actions swaps and evaluation performance with different quantities of $\epsilon$-actions and color jitter (both only applied for evaluation). All measurements use the final agent, trained on 200 million frames, for Atari Phoenix, averaged over 3 seeds. Action Gap is the average absolute Q-value difference between the highest two valued actions. % Actions Swap is the percentage of times the agent’s argmax action has changed from the last timestep. Policy churn is the percentage of states in which the agent’s argmax action has changed after a single gradient step. Color jitter applies a random 10% change to the brightness, saturation and hue of each frame. For associated errors with these values, please see Appendix E.3.
Category BTR w/o Munchausen w/o IQN w/o SN w/o Impala w/o Maxpool
Action Gap 0.282 0.055 0.180 0.274 0.215 0.264
% Action Swaps 36.6% 47.7% 42.2% 40.3% 41.1% 39.3%
Policy Churn 3.8% 11.0% 0.5% 3.3% 4.5% 4.2%
Score ColorJitter 212k 85k 110k 187k 19k 187k
Score $\epsilon=0.03$ 94k 42k 62k 75k 10k 86k
Score $\epsilon=0.01$ 194k 70k 110k 132k 13k 171k
Score $\epsilon=0$ 330k 184k 187k 296k 21k 406k
Figure 7: Plot showing % of dormant neurons (Sokar et al., 2023), SRank with $\delta=0.01$ (Kumar et al., 2021) and L2 norm of network weights, for details see Appendix E.3. Results are averaged over 3 seeds and 5 tasks (Atari-5). Shading shows 95% confidence intervals.
@@ 5.2 What are the effects of BTR’s components?
To help interpret the results in Section 5.1, Table 2 measures seven different attributes of the agent either during or after training: action gaps and action swaps (linked to causing approximation errors (Bellemare et al., 2016)); policy churn (which can cause excessive off-policyness and instability (Schaul et al., 2022)) and score with additional noise (indicating robustness of the policies).
@@
While it is clear that Impala strongly contributes to performance, we find that without BTR’s other components the learned policy is highly noisy and unstable. Table 2, demonstrates that without IQN and Munchausen the agent experiences very low action gaps (absolute Q-value difference between the highest two valued actions), causing the agent to swap its argmax action almost every other step. This is likely to result in approximation errors altering the policy and causing a high degree of off-policyness in the replay buffer. This is particularly detrimental in games requiring fine-grained control, such as Phoenix where the agent needs to narrowly dodge many projectiles, reflected in BTR’s performance without these components.
@@
Furthermore, we find that maxpooling produces a more robust policy. To test this, we evaluate the performance of BTR’s ablations when taking different quantities of $\epsilon$-actions and with altered observations and find maxpooling alleviates some of the performance loss (Table 2). Lastly, we find Munchausen and IQN to have a significant impact on Policy Churn (Schaul et al., 2022), with Munchausen reducing it by 6.4% and IQN increasing it by 3.3%. As a result, when these components are used together, they appear to reach a level of churn which does not harm learning and potentially provides some exploratory benefits.
Lastly, Figure 7 shows an analysis of the trained model weights across the Atari-5 benchmark. We find little difference between trained models other than when removing Impala, which decreases dormant neurons and increases the L2 Norm of different layers, which have been linked with plasticity loss (Lyle et al., 2024).
@@
6 Related Work
The most similar work to BTR, developing a computationally-limited non-distributed RL algorithm, is “Fast and Efficient Rainbow” (Schmidt & Schmied, 2021). They optimized Rainbow DQN to maximize performance for 10 million frames through parallelizing the environments and dropping C51 along with hyperparameter optimizations. This differs from our goals of producing an algorithm that scales across training regimes (up to 200 million frames) and domains (Atari, Procgen, Super Mario Galaxy, Mario Kart and Mortal Kombat), resulting in different design decisions.
For less computation-limited approaches, Ape-X (Horgan et al., 2018) was the first to explore highly distributed training, allowing agents to be trained on a billion frames in 120 hours through using $>$ 100 CPUs. Following this, Kapturowski et al. (2018) proposed R2D2 using a recurrent neural network, increasing sample efficiency but slowing down gradient updates by 38%. Agent57 (Badia et al., 2020a) was the first RL agent to achieve superhuman performance across 57 Atari games, though required 90 billion frames. MEME (Kapturowski et al., 2023), Agent57’s successor, focused on achieving superhuman performance within the standard 200 million frames limit, achieved by using a significantly higher replay ratio and larger network architecture. Most recently, Dreamer-v3 (Hafner et al., 2023) used a 200 million parameter model requiring over a week of training, achieving similar results as MEME. We detail some key differences between BTR, MEME and Dreamer-v3 in Table 3. While these approaches perform equally or better than BTR, all are inaccessible to smaller research labs or hobbyists due to their required computational resources and walltime. Therefore, while these algorithms have important research value demonstrating the possible performance of RL agents, performative algorithms with a lower cost of entry, like BTR, are necessary for RL to become widely applicable and accessible.
7 Conclusions
@@
We acknowledge there exists many more promising improvements we could not include in BTR, leaving room for more future work to create stronger integrated agents in a few years. For example, BTR does not add an explicit exploration component, resulting in it struggling in hard-exploration tasks such as Montezuma’s Revenge; therefore, mechanisms used in Never Give Up (Badia et al., 2020b), etc may prove useful. Section 5.1 found that the neural network’s core architecture, Impala, had the largest impact on performance, an area we believe is generally underappreciated in RL. Previous work (Kapturowski et al., 2018) has incorporated recurrent models enhancing performance, though we are uncertain how this can be incorporated into BTR without affecting its computational accessibility, a question which warrants future research.
Impact Statement
This paper presents work whose goal is to advance the field of Reinforcement Learning, particularly to improve accessibility to those with limited computational resources. As with any work increasing accessibility, this has potential for misuse by bad actors. However, we believe these concerns are offset by the field’s potential to tackle key societal problems.
Acknowledgments
This work was supported by the UK Research and Innovation (UKRI) Centre for Doctoral Training in Machine Intelligence for Nano-electronic Devices and Systems [EP/S024298/1] and the Engineering and Physical Sciences Research Council (EPSRC) ActivATOR project [EP/W017466/1]. The authors acknowledge the use of the IRIDIS X High Performance Computing Facility, and the Southampton-Wolfson AI Research Machine (SWARM) GPU cluster generously funded by the Wolfson Foundation, together with the associated support services at the University of Southampton in the completion of this work. The authors dedicate this work to the memory of George Morton-Fallows, whose passion for computer science inspired this research.